OpenStudy (anonymous):

So I have this function: \[f(x)=(1+x)^{-3}\] And I need to find the Maclaurin's serie. I have found the following: \[=f(0)+\frac{f'(0)}{1!}x+\frac{f''(0)x^2}{2!}+...\] \[=1-\frac{3x}{1!}+\frac{12x^2}{2!}-\frac{60x^3}{3!}+...\] How do I convert this to a Sum?

OpenStudy (solomonzelman):

I will rely on your result, \(\large\color{#000000 }{ \displaystyle \sum_{k=0}^{\infty} \frac{}{k!}}\) but up to here it is fairly obvious if you look at the denominators, and recall also that 0!=1.

OpenStudy (solomonzelman):

Something happened to the equation in your question, but I can still read it...

OpenStudy (solomonzelman):

wait, is there a reason you removed it? you think it's incorrect?

OpenStudy (anonymous):

I havent removed anything?

OpenStudy (solomonzelman):

oh you didn't remove it and my latex is just not readin

OpenStudy (anonymous):

I think it is bugging. But I wrote: So I have this function: \[f(x)=(1+x)^{-3}\] And I need to find the Maclaurin's serie. I have found the following: \[=f(0)+\frac{f'(0)}{1!}x+\frac{f''(0)x^2}{2!}+...\] \[=1-\frac{3x}{1!}+\frac{12x^2}{2!}-\frac{60x^3}{3!}+...\] How do I convert this to a Sum?

OpenStudy (anonymous):

Yea, so I am having trouble finding the nominator.

OpenStudy (solomonzelman):

I missed it, I will erase the mess I made.

OpenStudy (anonymous):

\[\sum_{k=0}^{\infty}(-1)^k \frac{ (k+1)(k+2) }{ 2 }\]

OpenStudy (anonymous):

Wouldnt that make sense?

OpenStudy (anonymous):

missed \[x^n\] at the end

OpenStudy (anonymous):

\[\sum_{k=0}^{\infty}(-1)^k \frac{ (k+1)(k+2) }{ 2 }x^k\]

OpenStudy (solomonzelman):

and without k! ?

OpenStudy (anonymous):

Yea, but maybe this isnt a maclaurins serie?

OpenStudy (solomonzelman):

I will see if I can redo it and rethink it...

OpenStudy (solomonzelman):

I am lagging I need more time because that; apologize.

OpenStudy (anonymous):

Dont worry

OpenStudy (solomonzelman):

I am lagging so bad I am sorry

OpenStudy (anonymous):

No worries, ill take a look at some other questions. Thanks for your help though :)

OpenStudy (solomonzelman):

\[\sum_{k=0}^{\infty}(-1)^k \frac{ (k+1)(k+2) }{ 2 }x^k\] was correct

OpenStudy (solomonzelman):

Sorry for wasting your time.... (a bit frusturating with this internet. I got to go offline)

OpenStudy (anonymous):

No worries, have a great day